Please show how to solve this problem

2) Bargaining Suppose Rick is looking for a new assistant for his upcoming adventure to harvest Mega Seeds. Rick estimates that he can earn S5,000 if he goes alone or $7,000 if he hires Morty. If Morty doesnt take the trip hes stuck at home watching TV with Jerry, which is completely worthless to him. 1. What is Ricks threat value? Mortys? 2. What might be the wage that Rick offers Morty if he uses the reasonable solution bargaining outcome? What is each persons payoff? 3. What is the least amount Morty would accept to go on the adventure? What is the most Rick would be willing to pay Morty? 4. Suppose Rick invents a robot that costs him S1,000 to build, but will earn him $7,000 on the trip. How does this change the reasonable solution outcome? Why?
Answer #1

1. Under Nash bargaining strategy,

threat value of a player is equivalent to the probable loss due to disagreement

Here threat value of Rick = loss of income due to disagreement with Morty = $7000-$5000=$2000

Morty's threat value = loss of income due to refusal to work with Rick = $0

2. Rick's earning if he hires Morty =$7000

Under reasonable agreement of Nash bargaining, each party equally shares the earning

Therefore, payoff's of Rick= $7000/2=$3500 and payoff's of Morty = $7000/2= $3500

The reasonable agreement is a win-win situation for both parties as no one win or lose from bargaining process.

3. The Rick's threat value or loss of income if he does not hire Morty = $2000

therefore, under Nash bargaining model, the least amount that Morty will accept to work with Rick = $7000*1/(1+2000)

                                                                                                                      = 3.49

The maximum amount that Rick would be willing to pay Morty = $2000-$3.49= $1996.51

4. Cost of robot = $1000

Earning of Rick if he works with the help of robot = $7000

Therefore, net earning of Rick if he use robot in trip = $7000-$1000=$6000

Under reasonable agreement of Nash bargaining,

Payoff of Rick = $6000/2= $3000 and payoff of Morty = $6000/2=$3000